### v1

# One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings

Abstract
Deploying language models often requires handling model size vs. performance trade-offs to satisfy downstream latency constraints while preserving the model’s usefulness. Model distillation is commonly employed to reduce model size while maintaining acceptable performance. However, distillation can be inefficient since it involves multiple training steps. In this work, we introduce ModularStarEncoder, a modular multi-exit encoder with 1B parameters, useful for multiple tasks within the scope of code retrieval. ModularStarEncoder is trained with a novel self-distillation mechanism that significantly improves lower-layer representations—allowing different portions of the model to be used while still maintaining a good trade-off in terms of performance. Our architecture focuses on enhancing text-to-code and code-to-code search by systematically capturing syntactic and semantic structures across multiple levels of representation. Specific encoder layers are targeted as exit heads, allowing higher layers to guide earlier layers during training. This self-distillation effect improves intermediate representations, increasing retrieval recall at no extra training cost. In addition to the multi-exit scheme, our approach integrates a repository-level contextual loss that maximally utilizes the training context window, further enhancing the learned representations. We also release a new dataset constructed via code translation, seamlessly expanding traditional text-to-code benchmarks with code-to-code pairs across diverse programming languages. Experimental results highlight the benefits of self-distillation through multi-exit supervision.
One Model to Train them All: Hierarchical Self-Distillation for Enhanced
Early Layer Embeddings
Andrea Gurioli1, Federico Pennino1, Joao Monteiro2 Maurizio Gabbrielli1 1University of Bologna, 2Autodesk Correspondence: andrea.gurioli5@unibo.it, joao.monteiro@autodesk.com
1 Introduction
Large language models (LLMs) have significantly impacted the field of natural language processing, demonstrating remarkable performance across various applications (Niu et al., 2023). However, the amount of computation required to operate state-of-the-art models poses significant challenges for the large-scale deployment of these models.
To mitigate these challenges, the research community has explored several model strategies to reduce the operational cost of LLMs without sacrificing their effectiveness. A prominent technique in model compression is quantization (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024), which involves the reduction of numerical precision in the model’s parameters. Quantization effectively decreases memory requirements and enhances inference speed, facilitating the deployment of large language models in resource-constrained environments. Concurrently, knowledge distillation has emerged as a powerful technique whereby a smaller “student” model is trained to emulate the behavior of a larger “teacher” model, as evidenced by works such as DistilBERT (Sanh et al., 2019) and TinyBERT (Jiao et al., 2019). Additionally, pruning methods selectively eliminate less influential weights or neurons, further reducing model complexity and aiming to preserve performance (Han et al., 2015).
Recent efforts have increasingly focused on developing efficient architectures requiring fewer parameters. Model families such as LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a paradigm shift towards smaller, more accessible architectures. These model families are deployed at various resolutions—ranging from lightweight variants optimized for heavily resource-constrained environments to larger versions that retain competitive performance.
Figure 1: Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a “layer loss,” contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
In parallel, advancements in dynamic inference strategies have introduced mechanisms that further optimize computational efficiency. Techniques like multi-exit networks enable early predictions at intermediate layers, reducing unnecessary computations. For instance, early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) dynamically balance computation and accuracy by allowing predictions before full model execution. Similarly, Matryoshka representation learning (Kusupati et al., 2022) extends this idea to embeddings, introducing a loss function that yields multi-granular representations. This approach allows downstream tasks to adjust computational complexity by pruning embedding dimensionality, further contributing to efficient model deployment.
Building on these principles, we propose ModularStarEncoder, a modular multi-exit encoder architecture that integrates a novel intra-model self-distillation mechanism. In our design, specific intermediate layers are supervised by both the primary task loss and auxiliary distillation losses on specific exit heads, encouraging lower layers to learn better representations by mimicking the outputs of higher layers. We apply a shared embedding head comprising a masked language modeling head and an in-context classification head across a chosen subset of layers. We then fine-tuned the model with different projection heads for each exit point. We reached state-of-the-art results on multiple retrieval tasks (such as code-to-code and text-to-code), fine-tuning one single modular model that can be sliced depending on the end-user computational constraints.
Our contributions are as follows:
  • We introduce a self-distillation framework that enables training multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability. We believe this approach can significantly affect LLM training pipelines that depend on multiple model distillations.
  • We train and release ModularStarEncoder, which consists of a pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned encoder: The former is a modular pre-trained encoder with up to 1 billion parameters and five exit points, allowing users to perform multiple exit fine-tuning depending on downstream tasks. The latter is a fine-tuned encoder for various retrieval tasks. We allow the user to choose either the entire model with 1 billion parameters or a model size that fits their memory and computational limitations.
  • We release SynthCode2Code2NL a new dataset [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCode2Code2NL constructed via code translation, expanding popular text-to-code datasets across diverse programming languages with code-to-code pairs. SynthCode2Code2NL comprises 1 071 367 triplets of natural language-code-code.
2 Methodology
2.1 Dataset
Table 1: SynthCode2Code2NL details: Average character count and sample size per language for the CodeSearchNet dataset and the synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1 071 367 180 - -
PHP 280 706 514 116 967 579
Python 274 454 474 117 374 518
Go 234 089 350 124 125 541
Java 282 118 505 116 098 707
C++ - - 141 956 938
Ruby - - 158 494 456
C - - 136 365 1029
JavaScript - - 159 988 557
Figure 2: Prompt provided to Qwen2.5-Coder-7B-Instruct for translating a given code snippet ( print("Hello World") in the example) from a source programming language (Python) to a target one (Rust).
Table 2: Hyperparameters for Architecture, Pre-training, and Fine-tuning
Architecture
Hyperparameter Value
Model size 1B parameters
Precision bfloat16
Hidden layers 36
Attention heads 16
Hidden dimensionality 1024
Positional encoding RoPE ($\theta=10^{-6}$)
Context length 2048
Attention mechanism Grouped-Query Attention
Attention pattern Bi-directional
Pre-training
Batch size 3.99M tokens
Pretraining steps 245 000
Pretraining Tokens 1T
Loss function MLM + In-Context loss
Multi-layer loss yes
Optimizer AdamW
Weight decay 1e-1
Initial learning rate 6.24e-4
Learning rate schedule Multi-step
Warmup steps 4000
Fine-tuning
Dataset size 635.404 samples
Fine-tuning steps 20 000
Loss function CLIP loss
Multi-layer loss yes
Batch size 2048
Learning rate 1.0e-5
Temperature parameter 10.0
Hardware (Pre-training + fine-tuning)
GPUs 512 NVIDIA Ampere (64GB)
Overall Training hours 450 000
In the pre-training phase, we leveraged The Stack V2 Lozhkov et al. (2024), a large open-source code dataset structured by repository.
For the fine-tuning stage, we created SynthCode2Code2NL, a dataset that supports text-to-code and code-to-code search. Using the popular CodeSearchNet Husain et al. (2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 2). During code translation, we choose the token with the highest probability as output (greedy search) to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCode2Code2NL contains 1 071 367 samples where, in the first code column, we directly processed code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The third column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCode2Code2NL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams. Table 1 shows the average number of characters per language in SynthCode2Code2NL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in Limitations. Appendix A provides examples of code translation.
Figure 3: On the left side the illustration of the in-context loss framework, where samples from different repositories are concatenated. Positive examples share the same repository context, while negative examples come from different repositories. On the right side, in-context loss framework pseudocode.
2.2 Architecture
We updated the first version of StarEncoder (Li et al., 2023) by enabling longer code snippets (up to 2 048 tokens as context length), increasing the model size from $\approx$125M to $\approx$1B parameters and utilizing state-of-the-art methodologies (Warner et al., 2024; Lozhkov et al., 2024) resulting in ModularStarEncoder.
We built ModularStarEncoder on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its size from 15B to 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. ModularStarEncoder relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1024 with an intermediate size of 12 288.
We followed Devlin et al. (2019) and replaced the causal self-attention in StarCoder-2 with bidirectional self-attention. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. Table 2 summarizes the architectural details.
2.3 Pre-training
We pre-trained ModularStarEncoder with a batch size of 3.99M tokens for 245 000 training steps, processing $\approx$1T tokens. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs using the Leonardo supercomputer (Turisini et al., 2023), requiring 450 000 GPU working hours.
To enable both token-level and snippet-level embeddings after pre-training, we employed a multi-objective pre-training strategy that combined two losses, as detailed in section 2.3.1 and section 2.3.2. The pre-training was performed on TheStackV2, whose context length analysis revealed an average of $\approx 630$ tokens per code snippet. As described in section 2.3.1, we concatenated multiple snippets to facilitate our multi-loss methodology, allowing our in-context classification loss to expand the average context window to $\approx 1300$ tokens, reaching the maximum context length 20% of the time.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to 1e-6, and a weight decay of 1e-1. We initialized the learning rate at 6.24e-4 and decreased it using a multi-step learning rate schduler Bi et al. (2024) with 4 000 warmup steps. The learning rate was reduced at 120 000, 185 000, 220 000, 230 000, and 240 000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. Table 2 summarizes the hyperparameters for architecture, pre-training, and fine-tuning.
2.3.1 Masked Language Modeling and In-Context Classification
The training objectives of BERT (Feng et al., 2020), specifically Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), have become a de facto standard. However, The NSP loss constrains the context window length to the sentence length, leading to too many padding tokens and redundant computation (Zeng et al., 2022), and has been shown to not yield significant benefits after fine-tuning (Warner et al., 2024; Aroca-Ouellette and Rudzicz, 2020). Given that the average number of tokens per data sample in Stack v2 is 630, a large context window of 2048 results in substantial padding, making long-context training inefficient. While Wang et al. (2023) demonstrated the advantages of training LLMs with multiple objectives, we revisited the NSP loss and introduced an in-context classification (ICC) objective. We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments. Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head. Following Zhang et al. (2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy Devlin et al. (2019). The secondary objective, $\mathcal{L}_{ICC}$, determines whether randomly concatenated inputs (separated by a $<SEP>$ token) originate from the same repository (see fig. 3). Each concatenated sample has a 50% probability of containing source code from different repositories. This approach increases input density—reducing padding by expanding the average input length from 630 to 1 300 tokens—and potentially enhances cross-language understanding. Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may improve inter-language generalization.
Table 3: Performance of different models on text-to-code with CodeSearchNet using codeXGLUE. We reported the results presented in codet5plus, unixcoder and modernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
CodeSearchNet
Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
ModularStarEncoder 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
Codet5+ 770M 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
Unixcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
ModernBERT-large - - - - - - - 59.5
2.3.2 Multi-layer Loss
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.3.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This approach allows for flexible model deployment, enabling adaptive layer pruning while maintaining performance trade-offs.
2.4 Fine-tuning
Following Su et al. (2023), we fine-tune a single model for both text-to-code and code-to-code retrieval using instruction prompting. The optimization objective combines CLIP loss (Radford et al., 2021) with a multi-layer loss (details in 2.3.2).
To enhance representation learning, we replace the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2 048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $51.0e-$ 5 as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
2.5 Evaluation
We evaluated ModularStarEncoder fine-tuned, on both text-to-code and code-to-code retrieval tasks using CodeXGLUE (Lu et al., 2021), which comprises several benchmarking datasets. For text-to-code retrieval, we employed the CodeSearchNet dataset, where the goal is to retrieve the most relevant code snippet given a natural language query. Specifically, the query corresponds to a documentation comment, and the model is tasked with ranking the correct code snippet among 999 distractor snippets Husain et al. (2019). This setup assesses the model’s ability to learn meaningful cross-modal representations between code and natural language.
For code-to-code retrieval, we relied on two datasets from CodeXGLUE: the Code Translation (CT) benchmark and POJ-104. The Code Translation dataset consists of semantically equivalent code snippets in different programming languages, and we framed the task as cross-language code retrieval rather than translation. In this setting, given a Java code snippet as a query, the model retrieves the corresponding C# implementation, testing its capability to capture cross-lingual semantic similarities between functionally equivalent programs.
In contrast, withPOJ-104 dataset, we want to evaluate the model on intra-language semantic search (POJ-104 contains only C++ snippets), where programs solve the same problem but with different implementations. This setup evaluates the model’s capacity to generalize across structural variations while preserving semantic equivalence.
Table 4: Performance of different models on Code Translation (CT) and POJ104 for code-to-code search with codeXGLUE dataset.
CT POJ104
Model MRR mAP
ModularStarEncoder 98.9 56.5
Codet5+ 110M-embedding 98.4 24.5
OpenAI text-embedding-3-large 98.8 82.9
Unixcoder 97.6 41.0
ModernBERT-large 93.1 27.3
3 Results and Discussion
Table 5: Performance comparison of ModularStarEncoder layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by Mean Reciprocal Rank (MRR). We refer to ModularStarEncoder, fine-tuned with multiple exit points simultaneously, as self-distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
CodeSearchNet
Model Size Ruby Javascript Go Python Java PHP avg. MRR
Layer-4 $\approx 160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Layer-4 (self-distilled) 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Layer-9 $\approx 300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Layer-9 (self-distilled) 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Layer-18 $\approx 550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Layer-18 (self-distilled) 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Layer-27 $\approx 800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Layer-27 (self-distilled) 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Layer-36 $\approx 1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Layer-36 (self-distilled) 73.5 72.6 80.5 91.4 76.9 82.7 79.6
3.1 Benchmarks
Table 3 presents the results for CodeSearchNet (t2c) task in terms of Mean Reciprocal Rank (MRR) for each single language, average NDCG and average MRR. Results for Unixcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, ModularStarEncoder achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (Wang et al., 2023) (770M), UniXcoder (Guo et al., 2022), and ModernBERT-large (Warner et al., 2024). The only encoder that surpasses ModularStarEncoder is OpenAI’s text-embedding-3-large.
Table 4 presents results from both POJ104 and CT datasets reported respectively in MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++# retrieval). ModularStarEncoder reaches the best performance among the tests. We decided to replicate the benchmarking for all models in a zero-shot setting for code-to-code tasks because our model does not integrate POJ104 and the code translation datasets in the training set.
Referring to Table 4, on the POJ104 dataset in zero-shot, ModularStarEncoder achieves an mAP of 0.57, which is state-of-art between open-sourced models, however it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
3.2 Ablation Study
We conducted an ablation study by fine-tuning singularly each exit point (also starting from ModularStarEncoder, pre-trained) and pruning the subsequent layers (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just one projection head on that layer). Finally, we compared the sliced models with the corresponding results (self-distilled) of the model fine-tuned with the multi-layer loss (ModularStarEncoder). ModularStarEncoder consistently outperforms the single-exit baseline, indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 5, where ModularStarEncoder, indicated as self-distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in self-distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 4 illustrates that ModularStarEncoder maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval accuracy.
4 Related work
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019); Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
(a) MRR
(b) Recall@1
Figure 4: Performance Comparison Across Layers: The graph illustrates the MRR and the Recall@1 for different layers, comparing baseline models and a self-distilled model.
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with codet5plus, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
When trying to achieve better performance, research has shifted toward models with a high number of parameters. While this trend appears effective from a performance perspective, end users may face computational or memory limitations as LLMs vary from millions to billions of parameters. Sanh et al. (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior. This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
Our work differs from previous work by adapting a modern architecture Lozhkov et al. (2024) to a code encoder-only based model and introducing a novel ’self-distillation’ mechanism. We replace the next sentence prediction loss with an in-context classification focused on the repository level and expand the context to 2048 tokens. Our novel self-distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
5 Conclusion
In this work, we introduced ModularStarEncoder, a modular multi-exit encoder architecture designed to improve efficiency and scalability in code retrieval tasks. By integrating an intra-model self-distillation mechanism, our approach enables multiple resolution models to be trained within a unified layer stack, reducing redundancy while maintaining high retrieval performance. Our evaluation on CodeXGLUE demonstrates that ModularStarEncoder achieves state-of-the-art results among open-source models, outperforming prior baselines across text-to-code and code-to-code retrieval tasks. Ablations further highlighted the benefits of self-distillation, showing that lower layers gain representational strength from deeper layers, leading to superior performance compared to single-exit models.
Beyond performance gains, ModularStarEncoder offers practical benefits by providing multiple exit points, allowing users to balance computational efficiency and accuracy based on resource constraints. The results suggest that self-distillation provides a promising direction for efficient large-scale encoders, reducing deployment costs without sacrificing effectiveness.
Finally, released in open-access our SynthCode2Code2NL and both pre-trained and fine-tuned ModularStarEncoder models.
Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches with smaller and embryonic models were the only ways to extrapolate the best hyperparameter setup. The best hyperparameters for smaller models can differ from those for larger ones; thus, we faced a limitation in finding an optimal training setup. Ablating both the in-context classification and the multi-layer loss in a real scenario was impossible as we depended on smaller models to understand their performances. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor undermines the possibility of replicating the experiments.
Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
External validity
When synthesizing the SynthCode2Code2NL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
Internal validity
The ablation study focused on fine-tuning the model with and without multi-layer loss. However, this comparison does not account for how the model behaves when starting from a model not pre-trained on multi-layer loss. Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
Appendix A Synthetic dataset
SynthCode2Code2NL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCode2Code2NL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 5, In Figure 6 and Figure 7 some examples of code translation are shown.
Figure 5: Example of a code snippet translation from Go to Ruby.
Figure 6: Example of a code snippet translation from Java to C.
Figure 7: Example of a code snippet translation from Python to PHP.

### v2

@@
-One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings
+MoSE: Hierarchical Self-Distillation Enhances Early Layer Embeddings
 Abstract
-Deploying language models often requires handling model size vs.
-performance trade-offs to satisfy downstream latency constraints while preserving the model’s usefulness.
-Model distillation is commonly employed to reduce model size while maintaining acceptable performance.
-However, distillation can be inefficient since it involves multiple training steps.
-In this work, we introduce ModularStarEncoder, a modular multi-exit encoder with 1B parameters, useful for multiple tasks within the scope of code retrieval.
-ModularStarEncoder is trained with a novel self-distillation mechanism that significantly improves lower-layer representations—allowing different portions of the model to be used while still maintaining a good trade-off in terms of performance.
-Our architecture focuses on enhancing text-to-code and code-to-code search by systematically capturing syntactic and semantic structures across multiple levels of representation.
-Specific encoder layers are targeted as exit heads, allowing higher layers to guide earlier layers during training.
-This self-distillation effect improves intermediate representations, increasing retrieval recall at no extra training cost.
-In addition to the multi-exit scheme, our approach integrates a repository-level contextual loss that maximally utilizes the training context window, further enhancing the learned representations.
-We also release a new dataset constructed via code translation, seamlessly expanding traditional text-to-code benchmarks with code-to-code pairs across diverse programming languages.
-Experimental results highlight the benefits of self-distillation through multi-exit supervision.
-One Model to Train them All: Hierarchical Self-Distillation for Enhanced
-Early Layer Embeddings
-Andrea Gurioli1, Federico Pennino1, Joao Monteiro2 Maurizio Gabbrielli1 1University of Bologna, 2Autodesk Correspondence: andrea.gurioli5@unibo.it, joao.monteiro@autodesk.com
+Deploying language models often requires navigating accuracy vs.
+performance trade-offs to meet latency constraints while preserving utility.
+Traditional model distillation reduces size but incurs substantial costs through training separate models.
+We introduce ModularStarEncoder (MoSE), a 1-billion-parameter multi-exit encoder for code retrieval and classification that employs a novel Self-Distillation mechanism.
+This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs.
+Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training—improving intermediate representations at minimal additional cost.
+We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization.
+Additionally, we release a new dataset created through code translation that extends text-to-code benchmarks with cross-language code-to-code pairs.
+Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
 1 Introduction
-Large language models (LLMs) have significantly impacted the field of natural language processing, demonstrating remarkable performance across various applications (Niu et al., 2023).
-However, the amount of computation required to operate state-of-the-art models poses significant challenges for the large-scale deployment of these models.
-To mitigate these challenges, the research community has explored several model strategies to reduce the operational cost of LLMs without sacrificing their effectiveness.
-A prominent technique in model compression is quantization (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024), which involves the reduction of numerical precision in the model’s parameters.
-Quantization effectively decreases memory requirements and enhances inference speed, facilitating the deployment of large language models in resource-constrained environments.
-Concurrently, knowledge distillation has emerged as a powerful technique whereby a smaller “student” model is trained to emulate the behavior of a larger “teacher” model, as evidenced by works such as DistilBERT (Sanh et al., 2019) and TinyBERT (Jiao et al., 2019).
-Additionally, pruning methods selectively eliminate less influential weights or neurons, further reducing model complexity and aiming to preserve performance (Han et al., 2015).
-Recent efforts have increasingly focused on developing efficient architectures requiring fewer parameters.
-Model families such as LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a paradigm shift towards smaller, more accessible architectures.
-These model families are deployed at various resolutions—ranging from lightweight variants optimized for heavily resource-constrained environments to larger versions that retain competitive performance.
-Figure 1: Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36).
-Each exit head predicts an output embedding and adds a “layer loss,” contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
-In parallel, advancements in dynamic inference strategies have introduced mechanisms that further optimize computational efficiency.
-Techniques like multi-exit networks enable early predictions at intermediate layers, reducing unnecessary computations.
-For instance, early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) dynamically balance computation and accuracy by allowing predictions before full model execution.
-Similarly, Matryoshka representation learning (Kusupati et al., 2022) extends this idea to embeddings, introducing a loss function that yields multi-granular representations.
-This approach allows downstream tasks to adjust computational complexity by pruning embedding dimensionality, further contributing to efficient model deployment.
-Building on these principles, we propose ModularStarEncoder, a modular multi-exit encoder architecture that integrates a novel intra-model self-distillation mechanism.
-In our design, specific intermediate layers are supervised by both the primary task loss and auxiliary distillation losses on specific exit heads, encouraging lower layers to learn better representations by mimicking the outputs of higher layers.
-We apply a shared embedding head comprising a masked language modeling head and an in-context classification head across a chosen subset of layers.
-We then fine-tuned the model with different projection heads for each exit point.
-We reached state-of-the-art results on multiple retrieval tasks (such as code-to-code and text-to-code), fine-tuning one single modular model that can be sliced depending on the end-user computational constraints.
+Large language models (LLMs) have significantly impacted natural language processing (Niu et al., 2023), but their computational demands pose substantial deployment challenges.
+The field has responded with various efficiency strategies: quantization techniques reduce numerical precision (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al., 2019; Jiao et al., 2019); and pruning methods eliminate less influential weights (Han et al., 2015).
+Concurrently, model families like LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
+Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks.
+Early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al., 2022) enables adjustment of computational complexity through embedding dimensionality pruning.
+These approaches maintain performance while reducing computational demands in resource-constrained environments.
+Furthermore, insights from the SimCLR framework (Chen et al., 2020) revealed that the final layer of a model does not necessarily produce the most semantically meaningful representations.
+Valeriani et al.
+(2023) analyzed the geometric evolution of hidden representations within large transformer models across layers, outlining that intermediate layers are identified as holding the most semantically rich representations for downstream tasks.
+[…]
+Figure 1: (a) Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36).
+Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
+(b) Computational cost (GFLOPs) vs.
+performances trade-off at different exit layers over CodeSearchNet dataset (text-to-code in avg.
+MRR) and POJ104 (code-to-code in mAP).
+Despite a reduction of approximately 90% floating point operations from layer 36 to layer 4, MRR performance only drops by 6.4% in absolute terms.
+For POJ104, our best results are observed in the initial layers.
+Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al., 2024) architecture, with 1 billion parameters that incorporate intra-model self-distillation.
+As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective.
+This approach encourages lower layers to learn better representations by mimicking higher-layer outputs.
+Figure 1(b) demonstrates the resulting accuracy-computation trade-off of encoders trained under this framework: despite a 90% reduction in floating point operations from layer 36 to layer 4, performance over text-to-code retrieval drops by only 6.4% in absolute terms, while the model maintains the best performances for code-to-code retrieval in very early layers.
+Notably, the exit point with the best performance isn’t necessarily the final layer, as noted in previous findings due to Chen et al.
+(2020); Valeriani et al.
+(2023).
+This offers both deployment flexibility and a natural model selection mechanism for specific tasks.
+Users can select which portion of the model to use during inference, starting from the first exit point at 160M parameters.
+By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands.
+This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al., 2023).
+We fine-tuned MoSE on SynthCoNL (cf.
+Sec.
+5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko & Roy, 2015) for code clone detection, achieving state-of-the-art results among open models while maintaining deployment modularity.
 Our contributions are as follows:
-  • We introduce a self-distillation framework that enables training multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability.
-We believe this approach can significantly affect LLM training pipelines that depend on multiple model distillations.
-  • We train and release ModularStarEncoder, which consists of a pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned encoder: The former is a modular pre-trained encoder with up to 1 billion parameters and five exit points, allowing users to perform multiple exit fine-tuning depending on downstream tasks.
-The latter is a fine-tuned encoder for various retrieval tasks.
-We allow the user to choose either the entire model with 1 billion parameters or a model size that fits their memory and computational limitations.
-  • We release SynthCode2Code2NL a new dataset [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCode2Code2NL constructed via code translation, expanding popular text-to-code datasets across diverse programming languages with code-to-code pairs.
-SynthCode2Code2NL comprises 1 071 367 triplets of natural language-code-code.
-2 Methodology
-2.1 Dataset
-Table 1: SynthCode2Code2NL details: Average character count and sample size per language for the CodeSearchNet dataset and the synthesized portion obtained through translation.
-Language CSN samples CSN avg. char Synth. samples Synth. avg. char
-English 1 071 367 180 - -
-PHP 280 706 514 116 967 579
-Python 274 454 474 117 374 518
-Go 234 089 350 124 125 541
-Java 282 118 505 116 098 707
-C++ - - 141 956 938
-Ruby - - 158 494 456
-C - - 136 365 1029
-JavaScript - - 159 988 557
+  1. 1.
+    A self-distillation methodology that trains multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability—an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
+  2. 2.
+    MoSE, pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
+  3. 3.
+    SynthCoNL [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
+Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated.
+Positive examples originate from the same repository context, whereas negative examples are sourced from different repositories.
+To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence.
+On the right side, you’ll find the pseudocode for the in-context loss framework.
+2 Self-Distillation and In-Context Classification
+As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al., 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability.
+Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
+We hypothesize that applying supervision at multiple depths, not just at the output—encourages more robust and semantically rich intermediate representations.
+To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al., 2023).
+While we retain the widely adopted Masked Language Modeling (MLM) loss Feng et al.
+(2020), we discard the traditional Next Sentence Prediction (NSP) objective.
+NSP has been shown to offer minimal benefit after fine-tuning (Liu et al., 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially for long-form code inputs.
+In its place, we introduce an In-Context Classification (ICC) loss (see section 2.1) that increases input density by reducing padding and fosters better chunk-level comprehension, essential for tasks like semantic retrieval or classification.
+By integrating these two complementary objectives, MLM and ICC, and distributing their supervision across multiple layers, forming the basis of our Self-Distillation mechanism (see section 2.2), our training approach improves internal representation capabilities and flexibility, ultimately yielding a model better suited for modular deployment and early exiting.
+2.1 Masked Language Modeling and In-Context Classification
+We revisited the NSP loss and introduced an In-Context Classification (ICC) objective.
+We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments.
+Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
+[…]
+Table 1: Performance—reported in MRR, and when specified in NDCG—of different models on text-to-code with CodeSearchNet using CodeXGLUE.
+We reported the results presented in CodeT5+, UniXcoder and ModernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
+MoSE surpasses all existing open-source models, achieving an improvement of +3.6% compared to the best-performing open model, CodeT5+.
+MoSE reduced the performance gap between open-source models and closed-source ones such as OpenAI text-embedding-3-large.
+CodeSearchNet
+Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
+MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
+CodeT5+ (770M) 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
+UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
+ModernBERT-large - - - - - - - 59.5
+OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
+2.2 Self-Distillation: Multi-Layer loss
+To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index.
+The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied.
+The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set.
+This Self-Distillation mechanism, shown in Figure 1, allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
+3 Pre-training
+We built MoSE on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model.
+We reduced its original size, obtaining a model of 1B parameters.
+Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads.
+MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
+To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2.
+Aiming for modularity, we also replaced sliding window attention with full attention.
+This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al., 2021).
+Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference.
+We have included and summarized additional details about model architecture in appendix D.
+3.1 Training Details
+We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al., 2024) dataset.
+We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$.
+We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al., 2024) with 4,000 warmup steps.
+The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001.
+We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU working hours on the Leonardo supercomputer (Turisini et al., 2023).
+4 Downstream tasks
+4.1 Finetuning for Retrieval, text-to-code and code-to-code search
+Using the SynthCoNL dataset detailed in section 5, we leveraged a instruction prompting approach Su et al.
+(2023) and trained a single model for both text-to-code and code-to-code retrieval tasks.
+The optimization objective uses CLIP loss (Radford et al., 2021) with the same multi-layer approach discussed in section 2.2, modified to enhance representation learning.
+Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36).
+We used a batch of 2,048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
+We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings.
+We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged.
+After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
+Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with the state-of-the-art.
+BigCloneBench
+Model Recall Precision F1
+MoSE (L4) 96.4 89.8 93.0
+MoSE (L9) 96.6 90.4 93.4
+MoSE (L18) 96.4 92.1 94.2
+MoSE (L27) 96.5 91.8 94.1
+MoSE (L36) 96.5 91.7 94.1
+UniXcoder 92.9 97.6 95.2
+CodeBERT 94.7 93.4 94.1
+CodeT5+ (770M) 96.7 93.5 95.1
+4.2 Finetuning for classification, code clone detection
+[…]
+5 SynthCoNL
+Prompt Translate this “‘print("Hello World")”’ from Python to Rust. Here is the translated code: Qwen2.5-Coder-7B-Instruct Output fn main() { println!("Hello World!"); }
 Figure 3: Prompt provided to Qwen2.5-Coder-7B-Instruct for translating a given code snippet ( print("Hello World") in the example) from a source programming language (Python) to a target one (Rust).
-Table 2: Hyperparameters for Architecture, Pre-training, and Fine-tuning
-Architecture
-Hyperparameter Value
-Model size 1B parameters
-Precision bfloat16
-Hidden layers 36
-Attention heads 16
-Hidden dimensionality 1024
-Positional encoding RoPE ($\theta=10^{-6}$)
-Context length 2048
-Attention mechanism Grouped-Query Attention
-Attention pattern Bi-directional
-Pre-training
-Batch size 3.99M tokens
-Pretraining steps 245 000
-Pretraining Tokens 1T
-Loss function MLM + In-Context loss
-Multi-layer loss yes
-Optimizer AdamW
-Weight decay 1e-1
-Initial learning rate 6.24e-4
-Learning rate schedule Multi-step
-Warmup steps 4000
-Fine-tuning
-Dataset size 635.404 samples
-Fine-tuning steps 20 000
-Loss function CLIP loss
-Multi-layer loss yes
-Batch size 2048
-Learning rate 1.0e-5
-Temperature parameter 10.0
-Hardware (Pre-training + fine-tuning)
-GPUs 512 NVIDIA Ampere (64GB)
-Overall Training hours 450 000
-In the pre-training phase, we leveraged The Stack V2 Lozhkov et al. (2024), a large open-source code dataset structured by repository.
-For the fine-tuning stage, we created SynthCode2Code2NL, a dataset that supports text-to-code and code-to-code search.
-Using the popular CodeSearchNet Husain et al.
-(2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
+We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages.
+Using the popular CodeSearchNet (Husain et al., 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
 To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript.
 We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig.
-2).
-During code translation, we choose the token with the highest probability as output (greedy search) to prevent semantic discrepancies.
+3).
+During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies.
 This process yielded pairs of code snippets in distinct languages tied to the same natural language description.
 As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages.
-SynthCode2Code2NL contains 1 071 367 samples where, in the first code column, we directly processed code snippets from CodeSearchNet, including Python, Java, PHP, and Go.
-The third column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets.
+SynthCoNL contains 1,071,367 triplets where, in the first code column, we directly sampled code snippets from CodeSearchNet, including Python, Java, PHP, and Go.
+The second code column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets.
 After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments.
 Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port).
-During the preprocessing phase of SynthCode2Code2NL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column.
+During the preprocessing phase of SynthCoNL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column.
 During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams.
-Table 1 shows the average number of characters per language in SynthCode2Code2NL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in Limitations.
-Appendix A provides examples of code translation.
-Figure 3: On the left side the illustration of the in-context loss framework, where samples from different repositories are concatenated.
-Positive examples share the same repository context, while negative examples come from different repositories.
-On the right side, in-context loss framework pseudocode.
-2.2 Architecture
-We updated the first version of StarEncoder (Li et al., 2023) by enabling longer code snippets (up to 2 048 tokens as context length), increasing the model size from $\approx$125M to $\approx$1B parameters and utilizing state-of-the-art methodologies (Warner et al., 2024; Lozhkov et al., 2024) resulting in ModularStarEncoder.
-We built ModularStarEncoder on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model.
-We reduced its size from 15B to 1B parameters.
-Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads.
-ModularStarEncoder relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1024 with an intermediate size of 12 288.
-We followed Devlin et al.
-(2019) and replaced the causal self-attention in StarCoder-2 with bidirectional self-attention.
-Aiming for modularity, we also replaced sliding window attention with full attention.
-This step was taken to avoid the receptive field phenomenon of sliding window mechanisms (Zhu et al., 2021).
-Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference.
-Table 2 summarizes the architectural details.
-2.3 Pre-training
-We pre-trained ModularStarEncoder with a batch size of 3.99M tokens for 245 000 training steps, processing $\approx$1T tokens.
-We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs using the Leonardo supercomputer (Turisini et al., 2023), requiring 450 000 GPU working hours.
-To enable both token-level and snippet-level embeddings after pre-training, we employed a multi-objective pre-training strategy that combined two losses, as detailed in section 2.3.1 and section 2.3.2.
-The pre-training was performed on TheStackV2, whose context length analysis revealed an average of $\approx 630$ tokens per code snippet.
-As described in section 2.3.1, we concatenated multiple snippets to facilitate our multi-loss methodology, allowing our in-context classification loss to expand the average context window to $\approx 1300$ tokens, reaching the maximum context length 20% of the time.
-We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to 1e-6, and a weight decay of 1e-1.
-We initialized the learning rate at 6.24e-4 and decreased it using a multi-step learning rate schduler Bi et al.
-(2024) with 4 000 warmup steps.
-The learning rate was reduced at 120 000, 185 000, 220 000, 230 000, and 240 000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001.
-Table 2 summarizes the hyperparameters for architecture, pre-training, and fine-tuning.
-2.3.1 Masked Language Modeling and In-Context Classification
-The training objectives of BERT (Feng et al., 2020), specifically Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), have become a de facto standard.
-However, The NSP loss constrains the context window length to the sentence length, leading to too many padding tokens and redundant computation (Zeng et al., 2022), and has been shown to not yield significant benefits after fine-tuning (Warner et al., 2024; Aroca-Ouellette and Rudzicz, 2020).
-Given that the average number of tokens per data sample in Stack v2 is 630, a large context window of 2048 results in substantial padding, making long-context training inefficient.
-While Wang et al.
-(2023) demonstrated the advantages of training LLMs with multiple objectives, we revisited the NSP loss and introduced an in-context classification (ICC) objective.
-We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments.
-Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
-In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head.
-Following Zhang et al.
-(2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy Devlin et al.
-(2019).
-The secondary objective, $\mathcal{L}_{ICC}$, determines whether randomly concatenated inputs (separated by a $<SEP>$ token) originate from the same repository (see fig.
-3).
-Each concatenated sample has a 50% probability of containing source code from different repositories.
-This approach increases input density—reducing padding by expanding the average input length from 630 to 1 300 tokens—and potentially enhances cross-language understanding.
-Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may improve inter-language generalization.
-Table 3: Performance of different models on text-to-code with CodeSearchNet using codeXGLUE.
-We reported the results presented in codet5plus, unixcoder and modernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
+6 Results and Discussion
+6.1 Evaluation
+Figure 4: Performance of different models on CT and POJ104 for code-to-code retrieval with CodeXGLUE dataset.
+MoSE has state-of-the-art results with the CT benchmark and outperforms all the open-source models over the POJ104 benchmark.
+Compared to closed-source solutions (OpenAItext-embedding-3-large), MoSE performs on par in the code translation task and strongly reduces the gap between open- and closed-source models with the POJ104 benchmark.
+We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al., 2021) group of benchmarks.
+On CodeSearchNet dataset, we assessed cross-modal (natural language to code) learning by retrieving target code from 999 distractors using natural language queries.
+For code-to-code retrieval, again against 999 distractors, we utilized CodeXGLUE’s Code Translation (CT) dataset to test cross-lingual retrieval (e.g., Java to C# snippets implementing the same functionality), and its POJ104 dataset (C++ snippets, semantically equivalent but syntactically different) for intra-lingual semantic search, evaluating generalization over structural differences while preserving semantics.
+In our retrieval performance evaluation, when comparing MoSE with other encoders, we included results from OpenAI’s text-embedding-3-large as a representative commercial, closed-source model.
+Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko & Roy, 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
+6.2 Benchmarks
+Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR.
+Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023).
+On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT-large.
+The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
+In Figure 4, we present results from both POJ104 and CT datasets reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval).
+MoSE reaches the best performance among the evaluated models.
+We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks as our model does not integrate POJ104 and the code translation datasets in the training set.
+We present a comprehensive and detailed analysis of code retrieval errors in appendix C.
+Figure 5: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval.
+* indicates p < 0.05 and ** indicates p < 0.001.
+Different layers, despite a common training objective, yield different similarity scores.
+Referring to Figure 4, on the POJ104 dataset in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is state-of-the-art among open-sourced models.
+However, it is significantly behind OpenAI text-embedding-3-large.
+We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
+We observe a marked difference in performance depending on whether the task is framed as text-to-code or code-to-code retrieval.
+As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
+This representation divergence suggests that the input modality influences where semantic alignment is most effectively captured within the model.
+This observation aligns with Valeriani et al.
+(2023) findings, underlining that semantic information is distributed unevenly across layers and is better expressed in the intermediate ones.
+MoSE leverages this property by supporting dynamic exit points, enabling users to balance computational cost and task performance based on where the most relevant representations emerge.
+To further investigate this phenomenon, we compared similarity scores across layers to assess how each layer represents code snippets differently.
+Figure 5 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings.
+A detailed analysis of these results is provided in appendix C.
+In Table 2, we report classification performance across different exit layers of MoSE for the BigCloneBench benchmark.
+Results show stable performance throughout layers, with F1-scores from Layer 18 onward (94.2–94.1) closely matching those of deeper exits.
+Earlier layers (e.g., Layers 4 and 9) exhibit slightly lower precision, reflecting the expected trade-off between early inference and representational depth.
+Our method remains competitive compared to state-of-the-art models.
+While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
+6.3 Ablation Study
+We conducted an ablation study for the Multi-Layer loss, by fine-tuning (starting from MoSE, pre-trained) on SynthCoNL, each exit point singularly (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just the projection head on that layer).
+Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss; all tests are conducted on the CodeSearchNet dataset.
+MoSE consistently outperforms the single-exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower-level layers benefit from training signals propagated from deeper layers.
+This behavior is highlighted in Table 3, where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently.
+This finding underscores a promising new direction in Self-Distillation for large-scale code and text models, enabling high performance even in more compact configurations.
+Moreover, Figure 6 illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
+Table 3: Performance comparison of MoSE layers and baseline fine-tuned models on the CodeSearchNet benchmark.
+The table displays the overall retrieval performance measured by the Mean Reciprocal Rank (MRR, higher is better).
+We refer to MoSE, fine-tuned with multiple exit points simultaneously, as Self-Distilled.
+The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
+MoSE fine-tuned with multi-layer loss (Self-Distillation) consistently outperforms the baselines, showing a more significant difference in the initial layers.
 CodeSearchNet
-Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
-ModularStarEncoder 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
-Codet5+ 770M 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
-OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
-Unixcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
-ModernBERT-large - - - - - - - 59.5
-2.3.2 Multi-layer Loss
-To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.3.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index.
-The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied.
-The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set.
-This approach allows for flexible model deployment, enabling adaptive layer pruning while maintaining performance trade-offs.
-2.4 Fine-tuning
-Following Su et al.
-(2023), we fine-tune a single model for both text-to-code and code-to-code retrieval using instruction prompting.
-The optimization objective combines CLIP loss (Radford et al., 2021) with a multi-layer loss (details in 2.3.2).
-To enhance representation learning, we replace the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36).
-We used a batch of 2 048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
-We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings.
-We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged.
-After conducting a grid search, we selected $51.0e-$ 5 as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
-2.5 Evaluation
-We evaluated ModularStarEncoder fine-tuned, on both text-to-code and code-to-code retrieval tasks using CodeXGLUE (Lu et al., 2021), which comprises several benchmarking datasets.
-For text-to-code retrieval, we employed the CodeSearchNet dataset, where the goal is to retrieve the most relevant code snippet given a natural language query.
-Specifically, the query corresponds to a documentation comment, and the model is tasked with ranking the correct code snippet among 999 distractor snippets Husain et al.
-(2019).
-This setup assesses the model’s ability to learn meaningful cross-modal representations between code and natural language.
-For code-to-code retrieval, we relied on two datasets from CodeXGLUE: the Code Translation (CT) benchmark and POJ-104.
-The Code Translation dataset consists of semantically equivalent code snippets in different programming languages, and we framed the task as cross-language code retrieval rather than translation.
-In this setting, given a Java code snippet as a query, the model retrieves the corresponding C# implementation, testing its capability to capture cross-lingual semantic similarities between functionally equivalent programs.
-In contrast, withPOJ-104 dataset, we want to evaluate the model on intra-language semantic search (POJ-104 contains only C++ snippets), where programs solve the same problem but with different implementations.
-This setup evaluates the model’s capacity to generalize across structural variations while preserving semantic equivalence.
-Table 4: Performance of different models on Code Translation (CT) and POJ104 for code-to-code search with codeXGLUE dataset.
-CT POJ104
-Model MRR mAP
-ModularStarEncoder 98.9 56.5
-Codet5+ 110M-embedding 98.4 24.5
-OpenAI text-embedding-3-large 98.8 82.9
-Unixcoder 97.6 41.0
-ModernBERT-large 93.1 27.3
-3 Results and Discussion
-Table 5: Performance comparison of ModularStarEncoder layers and baseline fine-tuned models on the CodeSearchNet benchmark.
-The table displays the overall retrieval performance measured by Mean Reciprocal Rank (MRR).
-We refer to ModularStarEncoder, fine-tuned with multiple exit points simultaneously, as self-distilled.
-The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
-CodeSearchNet
-Model Size Ruby Javascript Go Python Java PHP avg. MRR
-Layer-4 $\approx 160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
-Layer-4 (self-distilled) 62.2 64.7 74.8 88.1 71.4 78.0 73.2
-Layer-9 $\approx 300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
-Layer-9 (self-distilled) 67.6 69.4 78.9 90.2 75.5 82.3 77.3
-Layer-18 $\approx 550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
-Layer-18 (self-distilled) 74.1 74.0 82.5 92.5 78.7 84.5 81.0
-Layer-27 $\approx 800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
-Layer-27 (self-distilled) 73.2 73.3 81.7 92.1 77.8 83.8 80.3
-Layer-36 $\approx 1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
-Layer-36 (self-distilled) 73.5 72.6 80.5 91.4 76.9 82.7 79.6
-3.1 Benchmarks
-Table 3 presents the results for CodeSearchNet (t2c) task in terms of Mean Reciprocal Rank (MRR) for each single language, average NDCG and average MRR.
-Results for Unixcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023).
-On CodeSearchNet, ModularStarEncoder achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (Wang et al., 2023) (770M), UniXcoder (Guo et al., 2022), and ModernBERT-large (Warner et al., 2024).
-The only encoder that surpasses ModularStarEncoder is OpenAI’s text-embedding-3-large.
-Table 4 presents results from both POJ104 and CT datasets reported respectively in MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++# retrieval).
-ModularStarEncoder reaches the best performance among the tests.
-We decided to replicate the benchmarking for all models in a zero-shot setting for code-to-code tasks because our model does not integrate POJ104 and the code translation datasets in the training set.
-Referring to Table 4, on the POJ104 dataset in zero-shot, ModularStarEncoder achieves an mAP of 0.57, which is state-of-art between open-sourced models, however it is significantly behind OpenAI text-embedding-3-large.
-We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
-3.2 Ablation Study
-We conducted an ablation study by fine-tuning singularly each exit point (also starting from ModularStarEncoder, pre-trained) and pruning the subsequent layers (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just one projection head on that layer).
-Finally, we compared the sliced models with the corresponding results (self-distilled) of the model fine-tuned with the multi-layer loss (ModularStarEncoder).
-ModularStarEncoder consistently outperforms the single-exit baseline, indicating that lower-level layers benefit from training signals propagated from deeper layers.
-This behavior is highlighted in Table 5, where ModularStarEncoder, indicated as self-distilled, outperforms all the single exit baselines consistently.
-This finding underscores a promising new direction in self-distillation for large-scale code and text models, enabling high performance even in more compact configurations.
-Moreover, Figure 4 illustrates that ModularStarEncoder maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval accuracy.
+Model type Num. Layers Num. Params. Ruby Javascript Go Python Java PHP avg. MRR
+Base 4 $160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
+Self-Distilled 62.2 64.7 74.8 88.1 71.4 78.0 73.2
+Base 9 $300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
+Self-Distilled 67.6 69.4 78.9 90.2 75.5 82.3 77.3
+Base 18 $550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
+Self-Distilled 74.1 74.0 82.5 92.5 78.7 84.5 81.0
+Base 27 $800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
+Self-Distilled 73.2 73.3 81.7 92.1 77.8 83.8 80.3
+Base 36 $1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
+Self-Distilled 73.5 72.6 80.5 91.4 76.9 82.7 79.6
 7 Related work
+Figure 6: Performance Comparison between MoSE fine-tuned with Multi-layer loss (Self-Distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set.
+Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to 3.47% MRR and +4.36% Recall@1).
 Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks.
@@
 Lan et al.
-(2019); Liu et al.
+(2019) and Liu et al.
 (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization.
 More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
-(a) MRR
-(b) Recall@1
-Figure 4: Performance Comparison Across Layers: The graph illustrates the MRR and the Recall@1 for different layers, comparing baseline models and a self-distilled model.
 In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language.
@@
 Wang et al.
-(2023) expanded these findings with codet5plus, stressing how multiple losses that leverage code semantics impact the model pertaining.
+(2023) expanded these findings with CodeT5+, stressing how multiple losses that leverage code semantics impact the model pertaining.
 The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
@@
 This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
-Our work differs from previous work by adapting a modern architecture Lozhkov et al.
-(2024) to a code encoder-only based model and introducing a novel ’self-distillation’ mechanism.
-We replace the next sentence prediction loss with an in-context classification focused on the repository level and expand the context to 2048 tokens.
+Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al., 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism.
+We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens.
 Our novel Self-Distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
 8 Conclusion
-In this work, we introduced ModularStarEncoder, a modular multi-exit encoder architecture designed to improve efficiency and scalability in code retrieval tasks.
-By integrating an intra-model self-distillation mechanism, our approach enables multiple resolution models to be trained within a unified layer stack, reducing redundancy while maintaining high retrieval performance.
-Our evaluation on CodeXGLUE demonstrates that ModularStarEncoder achieves state-of-the-art results among open-source models, outperforming prior baselines across text-to-code and code-to-code retrieval tasks.
-Ablations further highlighted the benefits of self-distillation, showing that lower layers gain representational strength from deeper layers, leading to superior performance compared to single-exit models.
-Beyond performance gains, ModularStarEncoder offers practical benefits by providing multiple exit points, allowing users to balance computational efficiency and accuracy based on resource constraints.
-The results suggest that self-distillation provides a promising direction for efficient large-scale encoders, reducing deployment costs without sacrificing effectiveness.
-Finally, released in open-access our SynthCode2Code2NL and both pre-trained and fine-tuned ModularStarEncoder models.
+We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving state-of-the-art results on code understanding benchmarks.
+By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints.
+Our approach not only allows for modular use but also enhances the alignment between task requirements and representation depth.
+This improves task-specific layer specialization, providing the end-user with a more adaptable model.
+Future work will explore combining exit-points and further understanding task type and depth interactions.
+While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact.
+We discuss these limitations in detail in appendix A.
 Acknowledgments
 We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
-Limitations
+This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
+Appendix A Limitations
 Due to our dependence on multiple GPUs, we encountered significant computational constraints.
-Parameter grid searches with smaller and embryonic models were the only ways to extrapolate the best hyperparameter setup.
-The best hyperparameters for smaller models can differ from those for larger ones; thus, we faced a limitation in finding an optimal training setup.
-Ablating both the in-context classification and the multi-layer loss in a real scenario was impossible as we depended on smaller models to understand their performances.
-Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor undermines the possibility of replicating the experiments.
+Parameter grid searches using smaller and earlier models were the only methods to extrapolate the best hyperparameter setup.
+The optimal hyperparameters for smaller models can differ from those for larger ones; consequently, we encountered a limitation in discovering an effective training setup.
+Conducting an ablation study for the in-context classification loss was not practicable due to the high costs of the pre-training stage.
+Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor can also undermine the possibility of replicating the experiments.
 Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
 External validity
-When synthesizing the SynthCode2Code2NL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans.
+When synthesizing the SynthCoNL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans.
 We tested the model’s performance on standard benchmarks.
@@
 Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
-Appendix A Synthetic dataset
-SynthCode2Code2NL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al.
+Appendix B Synthetic dataset
+Table 4: SynthCoNL composition: Average character count and sample size per language for the CodeSearchNet seed dataset and synthesized portion obtained through translation.
+Language CSN samples CSN avg. char Synth. samples Synth. avg. char
+English 1,071,367 180 - -
+PHP 280,706 514 116,967 579
+Python 274,454 474 117,374 518
+Go 234,089 350 124,125 541
+Java 282,118 505 116,098 707
+C++ - - 141,956 938
+Ruby - - 158,494 456
+C - - 136,365 1,029
+JavaScript - - 159,988 557
+SynthCoNL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al.
 (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript).
@@
 Each dataset entry consists of a natural language description and two code snippets in different languages.
-SynthCode2Code2NL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java).
+SynthCoNL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java).
 In Figure 11, In Figure 12 and Figure 10 some examples of code translation are shown.
+Table 4 shows the average number of characters per language in SynthCoNL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in the Limitations section.
+Appendix C Failure Case Analysis in Code-to-Code Retrieval
+Figure 7: Retrieved incorrect samples from CodeT5+, UniXcoder, and the 18th layer of MoSE.
+All the models appear to be misled by structural similarities, such as loop functions and return values, highlighting a common superficial retrieval phase.
+Figure 8: Histograms of normalized similarity score between correct and incorrect retrieved code snippets.
+The first 5 distributions belong respectively to MoSE 4, 9, 18, 27, 36 layers’ exit points; CodeT5+ and UniXcoder are the last ones displayed.
+While our model tends to have different, distinct distributions regarding wrong and correct retrievals, both UniXcoder and CodeT5+ have more overlapping results.
+Table 5: Permutation tests between correct and incorrect similarity scores ($\alpha=0.05$). The table shows a big discrepancy between MoSE t-statistics and both UniXcoder and CodeT5+ results.
+Model t-statistic p-value
+MoSE (L4) 8.0 <0.001
+MoSE (L9) 7.3 <0.001
+MoSE (L18) 8.0 <0.001
+MoSE (L27) 7.8 <0.001
+MoSE (L36) 6.3 <0.001
+UniXcoder 2.2 0.04
+CodeT5+ 110M-embedding 2.7 0.009
+Figure 9: Heatmap of permutation tests (10,000 permutations) between MoSE similarity score results ($\alpha=0.05$).
+* with p<0.05 and ** with p<0.001.
+In the left heatmap, we present the differences in correctly retrieved snippets across various layers, while the right heatmap illustrates the incorrectly retrieved snippets.
+The heatmaps demonstrate that even layers trained to achieve similar representation objectives can yield different similarity scores.
+In this section, we analyzed where our model fails in code-to-code retrieval, relying upon gfg, a Python-to-Java dataset.
+We compared how our model gets misled by different code snippets with CodeT5+ and UniXcoder, and the confidence level with which the model identifies the occurrences.
+In fig.
+7 we highlight a common defect of the most used open code encoders.
+When retrieving code snippets focusing on semantic similarity, we emphasize how all the inspected models might get misled by structural factors such as loops, if statements, and return types.
+The retrieved results underline an overconsideration of structural features despite semantic differences.
+We then analyzed whether the model’s confidence might suggest uncertainty in retrieving incorrect snippets of code, indicating a statistically significant level of "awareness" in overconsidering misleading features.
+Thus, we compared correct and incorrect similarity distribution scores for retrieved code snippets.
+Figure 8 shows how MoSE, CodeT5+, and UniXcoder results differ when comparing correct and incorrect retrievals.
+In table 5, MoSE demonstrated a more significant t-statistic than CodeT5+ and UniXcoder.
+In a retrieval use case scenario, when the confidence level is insufficient, we suggest that thresholding the similarity score could be an option when deciding whether to provide the end user with a single result or a herd of results.
+In fig.
+9, we present the differences in similarity scores among the five exit points using a permutation test.
+It is important to note that, although the model is trained to maintain similar retrieval capabilities across these five exit points (as discussed in section 2.1), and while the incorrectly retrieved snippets might overlap between the layers, each exit point encodes code snippets differently.
+Based on these results, we strongly recommend conducting a grid search between the similarity score threshold and the exit point to utilize.
+Figure 10: Example of a code snippet translation from Python to PHP.
 Figure 11: Example of a code snippet translation from Go to Ruby.
 Figure 12: Example of a code snippet translation from Java to C.
-Figure 7: Example of a code snippet translation from Python to PHP.
+Appendix D Model card
+Table 6 provides a comprehensive overview of the hyperparameters, architectural choices, and training configurations used in the development of MoSE.
+This includes detailed settings for both the pre-training and fine-tuning stages, as well as hardware specifications.
+The information is provided to enhance reproducibility and improve understanding of the model’s design.
+Table 6: Details on hyperparameters, architecture, and training for MoSE are presented here, including insights from both pre-training and fine-tuning stages.
+Architecture Pre-training
+Hyperparameter Value Hyperparameter Value
+Model size 1B parameters Batch size 3.99M tokens
+Precision bfloat16 Pretraining steps 245,000
+Hidden layers 36 Pretraining tokens 1T
+Attention heads 16 Loss function MLM + In-Context loss
+Hidden dimensionality 1,024 Multi-layer loss Yes
+Positional encoding RoPE ($\theta=10^{-6}$) Optimizer AdamW
+Context length 2,048 Weight decay $1\mathrm{e}{-1}$
+Attention mechanism Grouped Query Attention Initial learning rate $6.24\mathrm{e}{-4}$
+Attention pattern Bi-directional Learning rate schedule Multi-step
+Warmup steps 4,000
+Fine-tuning Hardware (Pre-training + Fine-tuning)
+Hyperparameter Value Hyperparameter Value
+Dataset size 635,404 samples Pre-training GPUs 512 NVIDIA Ampere A100 (64GB)
+Fine-tuning steps 20,000 Fine-tuning GPUs 512 NVIDIA Ampere A100 (64GB)
+Loss function CLIP loss Benchmarking GPUs 1 NVIDIA Ampere A100 (64GB)
+Multi-layer loss Yes Overall training hours 450,000
+Batch size 2,048
+Learning rate $1.0\mathrm{e}{-5}$
+Temperature parameter 10.0

### v3

@@
 This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs.
-Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training—improving intermediate representations at minimal additional cost.
+Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training, thereby improving intermediate representations at minimal additional cost.
 We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization.
@@
 Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
+MoSEhttps://huggingface.co/modularStarEncoder
+SynthCoNLhttps://huggingface.co/datasets/modularStarEncoder/SynthCoNL
 1 Introduction
-Large language models (LLMs) have significantly impacted natural language processing (Niu et al., 2023), but their computational demands pose substantial deployment challenges.
-The field has responded with various efficiency strategies: quantization techniques reduce numerical precision (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al., 2019; Jiao et al., 2019); and pruning methods eliminate less influential weights (Han et al., 2015).
-Concurrently, model families like LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
+Large language models (LLMs) have significantly impacted natural language processing (Niu et al.
+2023), but their computational demands pose substantial deployment challenges.
+The field has responded with various strategies aimed at improving efficiency: quantization techniques reduce numerical precision (Jacob et al.
+2017; Lin et al.
+2023; Egiazarian et al.
+2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al.
+2019; Jiao et al.
+2019); and pruning methods eliminate less influential parameters (Han et al.
+2015).
+Concurrently, model families like LLaMA (Dubey et al.
+2024), Qwen (Hui et al.
+2024), Mistral (Jiang et al.
+2023), and SmolLM (Allal et al.
+2025) exemplify a shift toward more efficient architectures at varying parameter scales.
 Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks.
-Early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al., 2022) enables adjustment of computational complexity through embedding dimensionality pruning.
+Early-exit architectures such as BranchyNet (Teerapittayanon et al.
+2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al.
+2022) enables adjustment of computational complexity through embedding dimensionality pruning.
 These approaches maintain performance while reducing computational demands in resource-constrained environments.
-Furthermore, insights from the SimCLR framework (Chen et al., 2020) revealed that the final layer of a model does not necessarily produce the most semantically meaningful representations.
+Furthermore, insights from the SimCLR framework (Chen et al.
+2020) revealed that the final layer of a model does not necessarily produce the most useful representations.
 Valeriani et al.
@@
 […]
-Figure 1: (a) Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36).
+Figure 1: (a) Overview of our multi‐exit Self-Distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36).
 Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
@@
 For POJ104, our best results are observed in the initial layers.
-Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al., 2024) architecture, with 1 billion parameters that incorporate intra-model self-distillation.
+Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al.
+2024) architecture, with 1 billion parameters that incorporate intra-model Self-Distillation.
 As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective.
@@
 By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands.
-This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al., 2023).
+This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al.
+2023).
 We fine-tuned MoSE on SynthCoNL (cf.
 Sec.
-5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko & Roy, 2015) for code clone detection, achieving state-of-the-art results among open models while maintaining deployment modularity.
+5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko and Roy 2015) for code clone detection, achieving competitive results among open models while maintaining deployment modularity.
 Our contributions are as follows:
   1. 1.
-    A self-distillation methodology that trains multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability—an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
+    A Self-Distillation methodology that trains multiple models within a unified layer stack, reducing redundancy and improving scalability, an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
   2. 2.
-    MoSE, pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
+    MoSE, pre-trained and fine-tuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
   3. 3.
-    SynthCoNL [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
+    SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
 Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated.
@@
 To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence.
-On the right side, you’ll find the pseudocode for the in-context loss framework.
+On the right side, you’$l$ find the pseudocode for the in-context loss framework.
 2 Self-Distillation and In-Context Classification
-As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al., 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability.
+As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al.
+2025), we design our pre-training strategy to support modularity and ease downstream task adaptability.
 Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
-We hypothesize that applying supervision at multiple depths, not just at the output—encourages more robust and semantically rich intermediate representations.
-To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al., 2023).
-While we retain the widely adopted Masked Language Modeling (MLM) loss Feng et al.
-(2020), we discard the traditional Next Sentence Prediction (NSP) objective.
-NSP has been shown to offer minimal benefit after fine-tuning (Liu et al., 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially for long-form code inputs.
-In its place, we introduce an In-Context Classification (ICC) loss (see section 2.1) that increases input density by reducing padding and fosters better chunk-level comprehension, essential for tasks like semantic retrieval or classification.
+We hypothesize that applying supervision at multiple depths, not just at the output, encourages more robust and semantically rich intermediate representations.
+To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al.
+2023).
+While we retain the widely adopted Masked Language Modeling (MLM) loss (Feng et al.
+2020a), we discard the traditional Next Sentence Prediction (NSP) objective.
+NSP has been shown to offer minimal benefit after fine-tuning (Liu et al.
+2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially so for long-form code inputs.
+In its place, we introduce an In-Context Classification (ICC) loss (see Section 2.1) that increases input density, reducing padding and fostering better chunk-level comprehension, essential for downstream tasks such as semantic retrieval or classification.
 By integrating these two complementary objectives, MLM and ICC, and distributing their supervision across multiple layers, forming the basis of our Self-Distillation mechanism (see Section 2.2), our training approach improves internal representation capabilities and flexibility, ultimately yielding a model better suited for modular deployment and early exiting.
@@
 Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
-[…]
-Table 1: Performance—reported in MRR, and when specified in NDCG—of different models on text-to-code with CodeSearchNet using CodeXGLUE.
-We reported the results presented in CodeT5+, UniXcoder and ModernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
-MoSE surpasses all existing open-source models, achieving an improvement of +3.6% compared to the best-performing open model, CodeT5+.
-MoSE reduced the performance gap between open-source models and closed-source ones such as OpenAI text-embedding-3-large.
-CodeSearchNet
-Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
-MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
-CodeT5+ (770M) 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
-UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
-ModernBERT-large - - - - - - - 59.5
-OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
+In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head.
+Following Zhang et al.
+(2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy (Devlin et al.
+2019).
+The secondary objective, $\mathcal{L}_{ICC}$, classifies whether randomly concatenated inputs (separated by a $\text{SEP}]$ special token) originate from the same repository (see Figure 2).
+Each concatenated sample has a 50% probability of containing source code from different repositories.
+This approach increases input density, reducing padding by expanding the average input length from 630 to 1,300 tokens.
+Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may also improve inter-language understanding.
+CodeSearchNet CT POJ104
+Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG MRR mAP
+MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2 98.9 75.9
+CodeT5+ 78.0 71.3 92.7 75.8 76.2 70.1 77.4 - 98.4 24.5
+UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 - 97.6 41.0
+ModernBERT-large - - - - - - - 59.5 93.1 27.3
+OpenAI Embedding 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3 98.8 82.9
+Table 1: Performance of different models on text-to-code and code-to-code tasks using the CodeXGLUE benchmark suite, specifically in terms of MRR (on CodeSearchNet and CodeTranslation CT dataset), NDCG, and mAP (on POJ104).
+On CodeSearchNet, MoSE surpasses all existing open-source models, achieving an improvement of 3.6% compared to the best-performing open model, CodeT5+.
+Furthermore, MoSE significantly reduces the performance gap between open-source models and closed-source alternatives, such as OpenAI text-embedding-3-large referred to as OpenAI Embedding, and performs comparably on the CT dataset.
+MoSE demonstrates competitive results with the CT benchmark and also outperforms all open-source models on the POJ104 benchmark.
 2.2 Self-Distillation: Multi-Layer loss
 To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (Section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index.
-The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied.
-The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set.
-This Self-Distillation mechanism, shown in Figure 1, allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
+The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha\$where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied.
+The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36).
+We also train a “tiny” single-exit baseline to enable isolating the Self-Distillation effect.
+The Self-Distillation mechanism illustrated in Figure 1 allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
 3 Pre-training
-We built MoSE on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model.
-We reduced its original size, obtaining a model of 1B parameters.
-Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads.
-MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
+We built MoSE on top of StarCoder-2 (Lozhkov et al.
+2024), applying several modifications to the model.
+We reduced its original size, resulting in a 1B-parameter model.
+Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al.
+2023) with 16 attention heads and 4 key-value heads.
+MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al.
+2021) with a base period $\theta=10^{6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
 To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2.
 Aiming for modularity, we also replaced sliding window attention with full attention.
-This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al., 2021).
-Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference.
-We have included and summarized additional details about model architecture in appendix D.
+This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al.
+2021).
+Finally, our implementation integrates FlashAttention V2 (Dao 2023) for faster inference.
+We have included and summarized additional details about the model architecture in the Appendix.
 3.1 Training Details
-We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al., 2024) dataset.
+We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al.
+2024) dataset.
 We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$.
-We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al., 2024) with 4,000 warmup steps.
+We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al.
+2024) with 4,000 warmup steps.
 The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001.
-We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU working hours on the Leonardo supercomputer (Turisini et al., 2023).
+We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU hours.
 4 Downstream Tasks
-4.1 Finetuning for Retrieval, text-to-code and code-to-code search
-Using the SynthCoNL dataset detailed in section 5, we leveraged a instruction prompting approach Su et al.
-(2023) and trained a single model for both text-to-code and code-to-code retrieval tasks.
-The optimization objective uses CLIP loss (Radford et al., 2021) with the same multi-layer approach discussed in section 2.2, modified to enhance representation learning.
+4.1 Fine-tuning for Retrieval, text-to-code and code-to-code search
+Using the SynthCoNL dataset detailed in Section 5, we leveraged an instruction prompting approach (Su et al.
+2023) and trained a single model for both text-to-code and code-to-code retrieval tasks.
+The optimization objective uses CLIP-style loss (Radford et al.
+2021) with the same multi-layer approach discussed in Section 2.2, modified to enhance representation learning.
 Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36).
-We used a batch of 2,048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
+We used a batch of 2,048 examples, ensuring that text-to-code and code-to-code were equally distributed across the batch.
 We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings.
 We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged.
-After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
-Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with the state-of-the-art.
+After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the fine-tuning process, and set the temperature parameter at 10.0.
 BigCloneBench
@@
 CodeT5+ (770M) 96.7 93.5 95.1
-4.2 Finetuning for classification, code clone detection
-[…]
+Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with open models.
+4.2 Fine-tuning for classification, code clone detection
+In addition to fine-tuning the model for multiple retrieval tasks as discussed in Section 4.1, we also explored how our methodology performs in a classification setup.
+We tested how the Multi-Layer and the In-Context classification losses adapt to the code clone detection task by: (1) fine-tuning multiple classification heads, one for each of the five exit points (layers 4, 9, 18, 27, and 36), and (2) leveraging the model’s ability to understand relationships between code snippets by inputting pairs of (hypothetical clone) code segments and training it to classify whether they are clones.
+We followed the input pattern devised in Section 2.1, formatting the input as follows: $\text{SEP}]$ snippet-1 $\text{SEP}]$ snippet-2 $\text{CLS}]$, and then used the final $\text{CLS}]$ token representation as input to a classifier.
+We fine-tuned the model with a constant learning rate of $1\mathrm{e}{-5}$ with 2,000 warmup steps, and used a batch size of 64 elements and trained for 14,000 steps.
 5 SynthCoNL
@@
 We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages.
-Using the popular CodeSearchNet (Husain et al., 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
+Using the popular CodeSearchNet (Husain et al.
+2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
 To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript.
-We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig.
-3).
+We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see Figure 3).
 During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies.
@@
 6.1 Evaluation
-Figure 4: Performance of different models on CT and POJ104 for code-to-code retrieval with CodeXGLUE dataset.
-MoSE has state-of-the-art results with the CT benchmark and outperforms all the open-source models over the POJ104 benchmark.
-Compared to closed-source solutions (OpenAItext-embedding-3-large), MoSE performs on par in the code translation task and strongly reduces the gap between open- and closed-source models with the POJ104 benchmark.
-We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al., 2021) group of benchmarks.
+We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al. 2021) group of benchmarks.
 On CodeSearchNet dataset, we assessed cross-modal (natural language to code) learning by retrieving target code from 999 distractors using natural language queries.
@@
 In our retrieval performance evaluation, when comparing MoSE with other encoders, we included results from OpenAI’s text-embedding-3-large as a representative commercial, closed-source model.
-Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko & Roy, 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
+Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko and Roy 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
 6.2 Benchmarks
 Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR.
-Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023).
-On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT-large.
+Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al.
+2022; Warner et al.
+2024; Wang et al.
+2023).
+On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT‐large.
 The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
-In Figure 4, we present results from both POJ104 and CT datasets reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval).
+In Table 1, we also present results from both POJ104 and CT datasets (code-to-code) reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval).
 MoSE reaches the best performance among the evaluated models.
-We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks as our model does not integrate POJ104 and the code translation datasets in the training set.
-We present a comprehensive and detailed analysis of code retrieval errors in appendix C.
+We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks, as our model does not integrate POJ104 and the code translation datasets in the training set.
+We present a comprehensive and detailed analysis of code retrieval errors in the Appendix.
 Figure 4: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval.
-* indicates p < 0.05 and ** indicates p < 0.001.
+* indicates p << 0.05 and ** indicates p << 0.001.
 Different layers, despite a common training objective, yield different similarity scores.
-Referring to Figure 4, on the POJ104 dataset in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is state-of-the-art among open-sourced models.
+On the POJ104 dataset, in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is the highest among open-source models to our knowledge.
 However, it is significantly behind OpenAI text-embedding-3-large.
-We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
-We observe a marked difference in performance depending on whether the task is framed as text-to-code or code-to-code retrieval.
+We underscore that a direct comparison with OpenAI text‐embedding‐3‐large remains challenging because it is closed‐source, and details such as model size, training methodology, or potential test data leakage are undisclosed.
+We observe a difference in performance profiles depending on whether the task is framed as text-to-code or code-to-code retrieval.
 As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
@@
 Figure 4 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings.
-A detailed analysis of these results is provided in appendix C.
+A detailed analysis of these results is provided in the Appendix.
 In Table 2, we report classification performance across different exit layers of MoSE for the BigCloneBench benchmark.
@@
 Earlier layers (e.g., Layers 4 and 9) exhibit slightly lower precision, reflecting the expected trade-off between early inference and representational depth.
-Our method remains competitive compared to state-of-the-art models.
+Our method remains competitive compared to open models.
 While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
 6.3 Ablation Study
-We conducted an ablation study for the Multi-Layer loss, by fine-tuning (starting from MoSE, pre-trained) on SynthCoNL, each exit point singularly (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just the projection head on that layer).
-Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss; all tests are conducted on the CodeSearchNet dataset.
-MoSE consistently outperforms the single-exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower-level layers benefit from training signals propagated from deeper layers.
-This behavior is highlighted in Table 3, where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently.
-This finding underscores a promising new direction in Self-Distillation for large-scale code and text models, enabling high performance even in more compact configurations.
-Moreover, Figure 6 illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
-Table 3: Performance comparison of MoSE layers and baseline fine-tuned models on the CodeSearchNet benchmark.
-The table displays the overall retrieval performance measured by the Mean Reciprocal Rank (MRR, higher is better).
-We refer to MoSE, fine-tuned with multiple exit points simultaneously, as Self-Distilled.
-The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
-MoSE fine-tuned with multi-layer loss (Self-Distillation) consistently outperforms the baselines, showing a more significant difference in the initial layers.
-CodeSearchNet
-Model type Num. Layers Num. Params. Ruby Javascript Go Python Java PHP avg. MRR
-Base 4 $160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
-Self-Distilled 62.2 64.7 74.8 88.1 71.4 78.0 73.2
-Base 9 $300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
-Self-Distilled 67.6 69.4 78.9 90.2 75.5 82.3 77.3
-Base 18 $550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
-Self-Distilled 74.1 74.0 82.5 92.5 78.7 84.5 81.0
-Base 27 $800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
-Self-Distilled 73.2 73.3 81.7 92.1 77.8 83.8 80.3
-Base 36 $1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
-Self-Distilled 73.5 72.6 80.5 91.4 76.9 82.7 79.6
+Effectiveness of the Multi Layer Loss
+We conducted an ablation study for the multi-layer loss, by fine-tuning each exit point independently on SynthCoNL, starting from the pre-trained MoSE.
+For example, for the baseline on layer 18, we retain the first 18 layers and fine-tune the model using just the projection head on that layer.
+Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss.
+All tests are conducted on the CodeSearchNet dataset.
+MoSE consistently outperforms the single‐exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower‐level layers benefit from training signals propagated from deeper layers.
+This behavior is highlighted in Figure 5 where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently.
+This finding underscores a promising new direction in Self‐Distillation for large‐scale code and text models, enabling high performance even in more compact configurations.
+Moreover, Figure 5 also illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
+Effectiveness of the In-Context Classification Objective
+CoIR (NDCG@10)
+Model CSN‑CCR CoSQA CT‑DL CT‑Contest
+MoSE (ICC) 10.1 0.4 32.0 12.1
+MoSE (NSP) 6.2 0.4 31.0 12.0
+Table 3: Ablation study comparing pre-training objectives (In-Context Classification vs.
+Next Sentence Prediction), evaluated with NDCG@10 (%) on the CoIR suite.
+ICC demonstrates gains on the cross-context CSN‑CCR task.
+We conducted an ablation study, detailed in Table 3, to assess whether our In-Context Classification (ICC) loss promotes inter-language understanding.
+For this study, we pre-trained two small versions of the model (each with 30M parameters) on 3B tokens.
+We compared the ICC loss with the Next Sentence Prediction (NSP) loss, commonly used in many state-of-the-art encoders (Li et al.
+2023; Feng et al.
+2020a).
+To assess the performance of our models, we utilized the inter-domain section of the CoIR evaluation suite (Li et al.
+2025), which focuses on tasks across various languages and domains.
+Our results indicated that the In-Context Classification loss either outperformed or matched the effectiveness of the Next Sentence Prediction loss, while providing a more efficient and dense representation.
+These results suggest that ICC may serve as a stronger default pretraining signal than NSP for multilingual and cross-context information retrieval tasks.
 7 Related Work
 Figure 5: Performance Comparison between MoSE fine-tuned with multi-layer loss (self-distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set.
-Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to 3.47% MRR and +4.36% Recall@1).
-Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks.
-BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses.
-Lan et al.
-(2019) and Liu et al.
-(2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization.
-More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
+Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to +4.36% Recall@1).
+Since the introduction of ELMo (Peters et al.
+2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks.
+BERT (Feng et al.
+2020a) followed those findings, adapting the Transformer architecture (Vaswani et al.
+2017) to enable a bi-directional representation with two different training objectives, namely the masked language modeling and the next sentence prediction losses.
+(Lan et al.
+2019) and (Liu et al.
+2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization.
+More recently, modernBERT (Warner et al.
+2024) tied the gap between modern decoders (Jiang et al.
+2023; Hui et al.
+2024; Dubey et al.
+2024; Touvron et al.
+2023; Lozhkov et al.
+2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
 In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language.
 Feng et al.
-(2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT.
-This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020).
-GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection.
-UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information.
+(2020b) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT.
+This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al.
+2020).
+GraphCodeBERT enhanced codeBERT (Feng et al.
+2020b) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection.
+UniXcoder (Guo et al.
+2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information.
 Wang et al.
@@
 (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior.
-This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
-Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al., 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism.
+This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al.
+2025; Hui et al.
+2024), becoming a standard for obtaining high-performing smaller LLMs.
+Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al.
+2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism.
 We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens.
@@
 8 Conclusion
-We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving state-of-the-art results on code understanding benchmarks.
+We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving competitive results on code understanding benchmarks.
 By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints.
@@
 While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact.
-We discuss these limitations in detail in appendix A.
 Acknowledgments
@@
 This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
-Appendix A Limitations
-Due to our dependence on multiple GPUs, we encountered significant computational constraints.
-Parameter grid searches using smaller and earlier models were the only methods to extrapolate the best hyperparameter setup.
-The optimal hyperparameters for smaller models can differ from those for larger ones; consequently, we encountered a limitation in discovering an effective training setup.
-Conducting an ablation study for the in-context classification loss was not practicable due to the high costs of the pre-training stage.
-Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor can also undermine the possibility of replicating the experiments.
-Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
-External validity
-When synthesizing the SynthCoNL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans.
-We tested the model’s performance on standard benchmarks.
-However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
-Internal validity
-The ablation study focused on fine-tuning the model with and without multi-layer loss.
-However, this comparison does not account for how the model behaves when starting from a model not pre-trained on multi-layer loss.
-Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
-Appendix B Synthetic dataset
-Table 4: SynthCoNL composition: Average character count and sample size per language for the CodeSearchNet seed dataset and synthesized portion obtained through translation.
-Language CSN samples CSN avg. char Synth. samples Synth. avg. char
-English 1,071,367 180 - -
-PHP 280,706 514 116,967 579
-Python 274,454 474 117,374 518
-Go 234,089 350 124,125 541
-Java 282,118 505 116,098 707
-C++ - - 141,956 938
-Ruby - - 158,494 456
-C - - 136,365 1,029
-JavaScript - - 159,988 557
-SynthCoNL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al.
-(2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript).
-The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns.
-Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
-For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency.
-Each dataset entry consists of a natural language description and two code snippets in different languages.
-SynthCoNL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java).
-In Figure 11, In Figure 12 and Figure 10 some examples of code translation are shown.
-Table 4 shows the average number of characters per language in SynthCoNL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in the Limitations section.
-Appendix C Failure Case Analysis in Code-to-Code Retrieval
-Figure 7: Retrieved incorrect samples from CodeT5+, UniXcoder, and the 18th layer of MoSE.
-All the models appear to be misled by structural similarities, such as loop functions and return values, highlighting a common superficial retrieval phase.
-Figure 8: Histograms of normalized similarity score between correct and incorrect retrieved code snippets.
-The first 5 distributions belong respectively to MoSE 4, 9, 18, 27, 36 layers’ exit points; CodeT5+ and UniXcoder are the last ones displayed.
-While our model tends to have different, distinct distributions regarding wrong and correct retrievals, both UniXcoder and CodeT5+ have more overlapping results.
-Table 5: Permutation tests between correct and incorrect similarity scores ($\alpha=0.05$). The table shows a big discrepancy between MoSE t-statistics and both UniXcoder and CodeT5+ results.
-Model t-statistic p-value
-MoSE (L4) 8.0 <0.001
-MoSE (L9) 7.3 <0.001
-MoSE (L18) 8.0 <0.001
-MoSE (L27) 7.8 <0.001
-MoSE (L36) 6.3 <0.001
-UniXcoder 2.2 0.04
-CodeT5+ 110M-embedding 2.7 0.009
-Figure 9: Heatmap of permutation tests (10,000 permutations) between MoSE similarity score results ($\alpha=0.05$).
-* with p<0.05 and ** with p<0.001.
-In the left heatmap, we present the differences in correctly retrieved snippets across various layers, while the right heatmap illustrates the incorrectly retrieved snippets.
-The heatmaps demonstrate that even layers trained to achieve similar representation objectives can yield different similarity scores.
-In this section, we analyzed where our model fails in code-to-code retrieval, relying upon gfg, a Python-to-Java dataset.
-We compared how our model gets misled by different code snippets with CodeT5+ and UniXcoder, and the confidence level with which the model identifies the occurrences.
-In fig.
-7 we highlight a common defect of the most used open code encoders.
-When retrieving code snippets focusing on semantic similarity, we emphasize how all the inspected models might get misled by structural factors such as loops, if statements, and return types.
-The retrieved results underline an overconsideration of structural features despite semantic differences.
-We then analyzed whether the model’s confidence might suggest uncertainty in retrieving incorrect snippets of code, indicating a statistically significant level of "awareness" in overconsidering misleading features.
-Thus, we compared correct and incorrect similarity distribution scores for retrieved code snippets.
-Figure 8 shows how MoSE, CodeT5+, and UniXcoder results differ when comparing correct and incorrect retrievals.
-In table 5, MoSE demonstrated a more significant t-statistic than CodeT5+ and UniXcoder.
-In a retrieval use case scenario, when the confidence level is insufficient, we suggest that thresholding the similarity score could be an option when deciding whether to provide the end user with a single result or a herd of results.
-In fig.
-9, we present the differences in similarity scores among the five exit points using a permutation test.
-It is important to note that, although the model is trained to maintain similar retrieval capabilities across these five exit points (as discussed in section 2.1), and while the incorrectly retrieved snippets might overlap between the layers, each exit point encodes code snippets differently.
-Based on these results, we strongly recommend conducting a grid search between the similarity score threshold and the exit point to utilize.
-Figure 10: Example of a code snippet translation from Python to PHP.
-Figure 11: Example of a code snippet translation from Go to Ruby.
-Figure 12: Example of a code snippet translation from Java to C.
-Appendix D Model card
-Table 6 provides a comprehensive overview of the hyperparameters, architectural choices, and training configurations used in the development of MoSE.
-This includes detailed settings for both the pre-training and fine-tuning stages, as well as hardware specifications.
-The information is provided to enhance reproducibility and improve understanding of the model’s design.
-Table 6: Details on hyperparameters, architecture, and training for MoSE are presented here, including insights from both pre-training and fine-tuning stages.
-Architecture Pre-training
-Hyperparameter Value Hyperparameter Value
-Model size 1B parameters Batch size 3.99M tokens
-Precision bfloat16 Pretraining steps 245,000
-Hidden layers 36 Pretraining tokens 1T
-Attention heads 16 Loss function MLM + In-Context loss
-Hidden dimensionality 1,024 Multi-layer loss Yes
-Positional encoding RoPE ($\theta=10^{-6}$) Optimizer AdamW
-Context length 2,048 Weight decay $1\mathrm{e}{-1}$
-Attention mechanism Grouped Query Attention Initial learning rate $6.24\mathrm{e}{-4}$
-Attention pattern Bi-directional Learning rate schedule Multi-step
-Warmup steps 4,000
-Fine-tuning Hardware (Pre-training + Fine-tuning)
-Hyperparameter Value Hyperparameter Value
-Dataset size 635,404 samples Pre-training GPUs 512 NVIDIA Ampere A100 (64GB)
-Fine-tuning steps 20,000 Fine-tuning GPUs 512 NVIDIA Ampere A100 (64GB)
-Loss function CLIP loss Benchmarking GPUs 1 NVIDIA Ampere A100 (64GB)
-Multi-layer loss Yes Overall training hours 450,000
-Batch size 2,048
-Learning rate $1.0\mathrm{e}{-5}$
-Temperature parameter 10.0